OpenStudy (anonymous):

Please Help!!!!!!!! Solve the equations for the solutions, x and y . x + y = k 2x + 3y = k + 1 The lines intersect at __, __

OpenStudy (anonymous):

multiply the first equation by \(-3\) and get \[-3x-3y=3k\\ 2x+3y=k+1\]

OpenStudy (anonymous):

oops , make that \[-3x-3y=-3k\\ 2x+3y=k+1\]

OpenStudy (anonymous):

then add and get \(-x=-2k+1\) making \(x=2k-1\)